### v1

# Pretrained Reversible Generation as Unsupervised Visual Representation Learning

Abstract
Recent generative models based on score matching and flow matching have significantly advanced generation tasks, but their potential in discriminative tasks remains underexplored. Previous approaches, such as generative classifiers, have not fully leveraged the capabilities of these models for discriminative tasks due to their intricate designs. We propose Pretrained Reversible Generation (PRG), which extracts unsupervised representations by reversing the generative process of a pretrained continuous flow model. PRG effectively reuses unsupervised generative models, leveraging their high capacity to serve as robust and generalizable feature extractors for downstream tasks. Our method consistently outperforms prior approaches across multiple benchmarks, achieving state-of-the-art performance among generative model-based methods, including 78% top-1 accuracy on ImageNet. Extensive ablation studies further validate the effectiveness of our approach.
1 Introduction
Figure 1: Overview of the PRG as Unsupervised Visual Representation pipeline. Swiss-roll data is generated using $(x,y)=\left(t\cos(t),t\sin(t)\right)$ with $t\in[0,3\pi]$, where the color gradient from blue to red corresponds to $t$ increasing from 00 to $3\pi$. The yellow trajectory represents the generative process, with color intensity indicating direction. Green arrows depict the reverse generation process; Using the pretrained model, each step along the green path can be fine-tuned for downstream tasks.
Generative models formulated as continuous-time stochastic differential equations, including diffusion and flow models [lipman2023flow, liu2022flow, pmlr-v202-pooladian23a, Albergo2023], have shown remarkable proficiency in multi-modal content generation tasks [rombach2022high, Ho2022video, mittal2021symbolic], as well as in scientific modeling [abramson2024accurate]. By effectively learning high-dimensional distributions, these models can generate new, high-quality samples that resemble the original data.
While generative tasks necessitate models to reconstruct the entire data distribution, another category, known as discriminative tasks (e.g., image classification), involve learning representations that capture the essential underlying data structure and output discrete or continuous predictions based on learned representations. Previous works have attempted to enhance representation learning through various methods, with particular attention on unsupervised learning recently. These methods endeavor to construct generic proxy tasks that effectively acquire and leverage data for pretraining, thereby providing robust representations for different downstream tasks. [bengio2013representation, radford2015unsupervised, chen2016infogan, van2017neural, donahue2019large, chen2020generative].
Due to the extraordinary performance of diffusion models in generative tasks, recent studies have explored their potential for discriminative tasks [clark2024text, li2023your]. These studies often rely on extracting or combining representations from the internal layers of pretrained diffusion models, involving intricate and non-generalizable designs, yet usually lacking a clear explanation. Moreover, the performance of these methods exhibits significant room for improvement compared to current state-of-the-art techniques.
Inspired by the reversible property of continuous-time stochastic flow models, in this paper, we introduce a general, model-agnostic, and effective method to apply generative models to discriminative tasks by reversing the generative process into a meaningful inference process. This method does not require accessing any internal features of a pretrained generative model, thus avoiding the need for fixed architectural modules to capture data representations. Another advantage of this method is that it enables unsupervised representation learning, making it applicable to unsupervised learning tasks without additional cost.
We demonstrate that this method is effective across diverse generative models. Through theoretical analysis and empirical experiments, we explain why reversing a pretrained generative process can serve as an effective feature extractor for downstream tasks. Our experiments show that this method achieves competitive performance in both supervised and unsupervised image classification tasks.
Our contributions can be summarized as follows:
  • We propose a general, model-agnostic method for applying pretrained generative models to discriminative tasks.
  • We analyze the theoretical foundation of the method and provide empirical findings demonstrating its efficacy.
  • Extensive experiments demonstrate the robustness and generalizability of our method on several image classification datasets.
2 Preliminaries
We introduce the generative models and mathematical notations used in this work, focusing on training and inference within diffusion and flow models.
2.1 Continuous-time Stochastic Flow Models
We consider two continuous-time stochastic flow generative models: diffusion models and flow models. Let $t$ represent time, with the interval $[1,0]$ for the generation process and $[0,1]$ for the reverse of generation process. We denote the data and its transformations as $x_{t}$, where $x_{0}$ is the original data from the dataset and $x_{1}$ is the terminal state, achieved by reversing the generative process.
2.1.1 Diffusion Model
A diffusion model reverses a forward diffusion process governed by the following SDE:
$\mathrm{d}x=f(t,x_{t})\mathrm{d}t+g(t)\mathrm{d}w_{t},$ (1)
where $f(t,x_{t})$ is the drift term, $g(t)$ is the diffusion coefficient, and $w_{t}$ is the Wiener process.
For simplicity, we use an SDE with a linear drift:
$\mathrm{d}x=f(t)x_{t}\mathrm{d}t+g(t)\mathrm{d}w_{t}.$ (2)
The transition distribution of $x\in\mathbb{R}^{d}$ from time 00 to $t$ is given by:
$p(x_{t}|x_{0})\sim\mathcal{N}(x_{t}|\alpha_{t}x_{0},\sigma_{t}^{2}I),$ (3)
where $\alpha_{t}$ is a scaling factor and $\sigma_{t}^{2}$ is the variance.
The diffusion process is predefined, such as the Variance-Preserving SDE (VP-SDE) [song2020score] or the Generalized VP-SDE (GVP) [Albergo2023]. Full definitions are provided in the Appendix A. The marginal distribution $p(x_{t})$ of the diffusion process is fixed if the data distribution $p(x_{0})$ is stationary. Using the Fokker-Planck equation, an equivalent ODE can be derived with the same marginal distribution, known as the Probability Flow ODE [song2020score]:
$\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=v(x_{t})=f(t)x_{t}-\frac{1}{2}g^{2}(t)\nabla_{x_{t}}\log p(x_{t}).$ (4)
In practice, the velocity model $v(x_{t})$ or score function $\nabla_{x_{t}}\log p(x_{t})$ in Eq. 4 is parameterized as $v_{\theta}$ or $s_{\theta}$ using a neural network, such as in Diffusion Transformers [peebles2023scalable, ma2024sit].
2.1.2 Flow Model
A flow model defines a deterministic transformation from $x_{0}$ to $x_{1}$ via a parameterized velocity function $v(x_{t})$, governed by the ODE: $\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=v(x_{t})$. The marginal probability distribution $p(x_{t})$ evolves through the vector field $v(x_{t})$, following the continuity equation:
$\frac{\partial p(x_{t})}{\partial t}+\nabla_{x}\cdot(p(x_{t})v(x_{t}))=0,$ (5)
which transports $p(x_{0})$ to $p(x_{1})$ along the flow path.
For the Independent Conditional Flow Matching (ICFM) models [tong2024improving] considered in this paper, the flow path is predefined as $p(x_{t}|x_{0},x_{1})=\mathcal{N}(x_{t}|tx_{1}+(1-t)x_{0},\sigma^{2}I)$, with the velocity function $v(x_{t}|x_{0},x_{1})=x_{1}-x_{0}$. We also studied Optimal Transport Conditional Flow Matching (OTCFM) models [tong2024improving]. The marginal distribution $p(x_{t})$ remains stationary if both $p(x_{0})$ and $p(x_{1})$ are stationary.
2.2 Generative Model Training and Inference
Both the diffusion and flow models can be parameterized by $v_{\theta}(x_{t})$ and pretrained using the Conditional Flow Matching loss, denoted $\mathcal{L}_{\text{FM}}$ [lipman2022flow, tong2024improving]. In the vanilla case, the weighting factor $\lambda_{\mathrm{FM}}$ equals $1$. For diffusion models, score matching is also applicable, with $\lambda_{\mathrm{SM}}(t)=\sigma_{t}^{2}$ in the vanilla case.
$\mathcal{L}_{\text{FM}}$ $=\frac{1}{2}\mathbb{E}_{p(x_{t})}\left[\lambda_{\mathrm{FM}}(t)\|v_{\theta}(x_{t})-v(x_{t})\|^{2}\right]\mathrm{d}t,$ (6)
$\mathcal{L}_{\text{SM}}$ $=\frac{1}{2}\mathbb{E}_{p(x_{t})}\left[\lambda_{\mathrm{SM}}(t)\|s_{\theta}(x_{t})-\nabla_{x_{t}}\log p(x_{t})\|^{2}\right]\mathrm{d}t.$
Once trained, starting with $x_{1}$ sampled from a standard Gaussian, we can recover $x_{0}$ from $x_{1}$ or infer the latent variable $x_{1}$ from $x_{0}$ by reversing the generative process.
3 Related Work
Generative versus Discriminative Classifiers
Generative classifiers is first proposed in [li2023your, chen2023robust, clark2024text, chendiffusion, zimmermann2021score, pmlr-v162-nie22a] model the conditional distribution $p(x|y)$ and use Bayes’ rule to compute the probability $p(y|x)=\frac{p(x|y)p(y)}{p(x)}$. These classifiers are typically trained by maximizing the conditional log-likelihood: $\max_{\theta}\mathbb{E}_{(x,y)\sim p(x,y)}\log p(x|y)$.
When employing conditional diffusion models, $p(x|y)$ can be computed with the instantaneous change-of-variables formula [Chen2018, Grathwohl2018] or approximated via ELBO [sohl2015deep, ho2020denoising]. duvenaud2020your advocate an energy-based model to enhance calibration, robustness, and out-of-distribution detection. yang2022your propose HybViT, which combines diffusion models and discriminative classifiers to predict $p(y|x)$. zimmermann2021score introduce score-based generative classifiers but find them vulnerable to adversarial attacks. clark2024text and li2023your explore zero-shot generative classifiers using pretrained models like Stable Diffusion [rombach2022high] to denoise images based on text labels.
In contrast, our approach fine-tunes a classifier on top of the generative model, making it a discriminative classifier that directly estimates $p(y|x)$. Discriminative classifiers are often preferred for classification tasks [Vapnik1995, AndrewNIPS2001]. fetaya2019understanding suggest a trade-off between likelihood-based modeling and classification accuracy, implying that improving generation may hinder classification for generative classifier. However, our two-stage training scheme shows that enhancing the generative model in the pretraining stage can improve discriminative performance during the finetuning stage. This suggests a complementary relationship between likelihood-based training and classification accuracy, indicating that discriminative models can benefit from generative pretraining. Thus, we demonstrate a viable approach to leveraging generative models for discriminative tasks.
Representation Learning via Denoising Autoencoders
Generative models have been widely explored for visual representation learning [kingma2013auto, van2017neural, goodfellow2020generative]. Denoising autoencoders (DAEs) [vincent2008extracting, vincent2010stacked, bengio2013generalized, geras2014scheduled] learn robust representations by reconstructing corrupted input data, which can be interpreted through a manifold learning perspective [vincent2011]. Our method extends vincent2010stacked by leveraging reversible flow-based generative models like diffusion models.
Diffusion models [sohl2015deep, ho2020denoising, song2021maximum] have shown promise in generation, yet their potential for feature extraction is underexplored. chen2020generative pretrain a Transformer for autoregressive pixel prediction, demonstrating competitive results in representation learning. Their two-stage approach—pretraining a generative model followed by full model finetuning—bears similarity to ours. Additionally, their downsampling of images to low resolution mimics the effect of adding noise, analogous to denoising. However, their model and autoregressive generation scheme are unsuitable for reverse generation and lack efficiency. he2022masked pretrain a masked autoencoder, effectively performing denoising by reconstructing incomplete data. baranchuk2021label show diffusion models improve semantic segmentation, particularly with limited labeled data. sinha2021dc propose Diffusion-Decoding with contrastive learning to enhance representation quality, while mittal2023diffusion introduce a Diffusion-based Representation Learning (DRL) framework, leveraging denoising score matching.
Recently, xiang2023denoising investigate using denoising diffusion autoencoders for discriminative tasks via generative pre-training. Our method differs by using the reverse generative process to extract features, offering a more general and model-agnostic approach.
4 Method
To address the issues discussed in Section 3, We propose an approach follows the conventional two-stage training scheme commonly used in generative modeling applications. In the first stage, we pretrain diffusion or flow models in an unsupervised manner. In the second stage, we fine-tune the models for discriminative tasks by conducting inference in the reverse direction of the generative process.
4.1 Pretraining Mechanism
In the traditional autoencoder paradigm, representation learning aims to learn a network that maps the input data $X$ into a useful representation $Z$ for downstream tasks. We parameterize the encoder $p_{\theta}(z|x)$ with parameters $\theta$. A good representation should retain sufficient information about the input $X$. In information-theoretic terms [linsker1988application, vincent2010stacked], this corresponds to maximizing the mutual information $\mathcal{I}(X,Z)$ between the input random variable $X$ and its representation $Z$. Following the reasoning of vincent2010stacked, we show that pretraining diffusion or flow models maximizes the mutual information. Specifically, we aim to find the optimal $\theta^{*}$ that maximizes $\mathcal{I}(X,Z)$:
$\theta^{*}$ $=\arg\max_{\theta}\mathcal{I}(X,Z)$ (7)
$=\arg\max_{\theta}[\mathcal{H}(X)-\mathcal{H}(X|Z)]$
$=\arg\max_{\theta}[-\mathcal{H}(X|Z)]$
$=\arg\max_{\theta}\mathbb{E}_{p(z,x)}[\log p(x|z)].$
Suppose we have a decoder $\theta^{\prime}$, which approximates $p_{\theta^{\prime}}(x|z)$ to recover data $x$ from the latent variable $z$. By the non-negativity of the KL divergence, we have:
$\mathbb{D}_{\mathrm{KL}}[p(x|z)||p_{\theta^{\prime}}(x|z)]\geq 0.$ (8)
$\mathbb{E}_{p(z,x)}[\log p(x|z)]\geq\mathbb{E}_{p(z,x)}[\log p_{\theta^{\prime}}(x|z)].$ (9)
Thus, the right term serves as a lower bound on $\mathcal{I}(X,Z)$.
In our setting, we consider deterministic mappings for both the encoder and decoder, which are invertible. Let us denote the mapping $z=F_{\theta}(x)$, which implies $p_{\theta}(z|x)=\delta(z-F_{\theta}(x))$, which is Dirac delta function. Under these assumptions, we can rewrite the RHS of Eq. 9 as:
$\mathbb{E}_{p(z,x)}{[\log{p_{\theta^{\prime}}(x|z)}]}$ $=\mathbb{E}_{p(x)}{[\log{p_{\theta^{\prime}}(x|z=F_{\theta}(x))}]}$ (10)
$=-\mathbb{D}_{\mathrm{KL}}(p(x)||p_{\theta^{\prime}}(x|z=F_{\theta}(x)))$
$-\mathcal{H}(X),$
where $p(x)$ is an unknown data distribution that we can sample from. Eq.10 is for maximizing the log likelihood of reconstructing data $X$, or equivalently, minimizing KL divergence between the data and the generated distribution.
In continuous-time stochastic flow generative models, the inference and generation process are reversible via a same model, which means the decoder parameter $\theta^{\prime}$ and encoder parameter $\theta$ are shared parameters. However, we still use these two notation in the later part of this paper for distinguishing the inference and generation process through ODE solver with $\theta$ and $\theta^{\prime}$ respectively. For preventing any confusion, we denote $x_{0}$ being data variable $x$ amd $x_{1}$ being the inferred latent variable $z$. In practice, training generative models using score matching or flow matching objectives does not directly maximize the likelihood of reconstructing $x_{0}$ from $x_{1}$. However, as shown in previous works [song2021maximum, lu2022maximum, zheng2023improved], these methods provide a lower bound on the likelihood. Specifically, they establish that:
$\mathbb{D}_{\mathrm{KL}}(p(x_{0})||p_{\theta^{\prime}}(x_{0}))=\mathbb{D}_{\mathrm{KL}}(p(x_{1})||p_{\theta^{\prime}}(x_{1}))+\mathcal{L}_{\text{ODE}}$ (11)
$\leq\mathbb{D}_{\mathrm{KL}}(p(x_{1})||p_{\theta^{\prime}}(x_{1}))+\sqrt{\mathcal{L}_{\text{SM}}}\sqrt{\mathcal{L}_{\text{Fisher}}}.$
The term $\mathbb{D}_{\mathrm{KL}}(p(x_{1})||p_{\theta^{\prime}}(x_{1}))\approx 0$ because both $p(x_{1})$ and $p_{\theta^{\prime}}(x_{1})$ are approximately the same Gaussian distribution. The losses $\mathcal{L}_{\text{ODE}}$, $\mathcal{L}_{\text{SM}}$ and $\mathcal{L}_{\text{Fisher}}$ are defined as:
$\mathcal{L}_{\text{ODE}}$ $=\frac{1}{2}\mathbb{E}_{p(x_{t})}[g^{2}(t)(s_{\theta}(x)-\nabla\log p(x_{t}))^{T}$ (12)
$(\nabla\log p_{\theta^{\prime}}(x_{t})-\nabla\log p(x_{t}))]\mathrm{d}t$
$\mathcal{L}_{\text{SM}}$ $=\frac{1}{2}\mathbb{E}_{p(x_{t})}\left[g^{2}(t)\|s_{\theta}(x)-\nabla\log p(x_{t})\|^{2}\right]\mathrm{d}t$
$\mathcal{L}_{\text{Fisher}}$ $=\frac{1}{2}\mathbb{E}_{p(x_{t})}\left[g^{2}(t)\|\nabla\log p_{\theta^{\prime}}(x_{t})-\nabla\log p(x_{t})\|^{2}\right]\mathrm{d}t,$
where $g(t)$ is a special weighting function, $s_{\theta}(x_{t})$ is the score function parameterized by $\theta$, and $p(x_{t})$ is the data distribution at time $t$ along forward diffusion process.
From Eq. 11, we observe that using score matching objectives during training can help increase the data likelihood. Flow matching differs from score matching by employing a slightly different weighting factor in the loss function, so it can also be used to increase the likelihood, as shown in Eq. 6. Therefore, pretraining diffusion or flow models via score matching or flow matching can be viewed as maximizing the mutual information between the data and its representation encoded through the reverse generative process. This relationship between the mutual information and the training loss is guaranteed by the Eq. 11. This process transforms the data distribution into a Gaussian distribution while preserving as much information as possible.
4.2 Finetuning by Reversing Generation
Due to the invertibility of the generative process (Eq. 4), the latent variable $x_{1}$ corresponding to a data point $x_{0}$ can be obtained by reversing the generative process. We denote $x_{t}=F_{\theta}(x_{0})$ as the extracted features for downstream discriminative tasks, such as classification or regression.
This paper focuses on supervised learning for downstream tasks. Suppose we have a labeled dataset $\mathcal{D}_{\text{finetune}}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ is a data point and $y_{i}$ is its label, potentially differing from the pretraining dataset $\mathcal{D}_{\text{pretrain}}$. We introduce a classifier $p_{\phi}(y\mid z)$, parameterized by $\phi$, trained by minimizing the cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{N}\log p_{\phi}(y_{i}\mid F_{\theta}(x_{i})).$ (13)
Finetuning Strategy
Generative features may not be optimal for discriminative tasks, as reconstruction requires more details than classification. Hence, fine-tuning is necessary to adapt the model for downstream tasks.
A key question is whether the flow model should also be fine-tuned alongside the classifier. Due to the flow model’s large capacity, reusing the pretrained model seems preferable. Initially, we explored freezing the flow model while finetuning only the classifier. However, this approach underperformed, indicating that fine-tuning both the flow model and the classifier is necessary. Otherwise, a large classifier with extra parameters would be required, reducing efficiency. To fine-tune the entire model, we leverage neural ODEs [Chen2018] for efficient gradient computation, optimizing both $\theta$ and $\phi$ via backpropagation.
We observe that the generative features are often redundant for discriminative tasks and are pruned during fine-tuning, as we will empirically demonstrate in the experiments section. Additionally, we assess the robustness and generalization of these features against out-of-distribution samples through further experiments.
Classifier Design
We found that the features $x_{t}$ are highly linearly separable, allowing a simple neural network to perform well without compromising accuracy. In practice, we also apply label smoothing [muller2019does] to enhance generalization. In this paper, we use a two-layer MLP with tanh activations as an example. Notably, increasing the classifier’s size did not significantly improve performance, suggesting that the features extracted by the reversed generative process are already highly informative.
Figure 2: CIFAR-10 Accuracy (y-axis) after fine-tuning checkpoints from different pretraining stages (x-axis).
Figure 3: Evaluation of the indicator (y-axis) $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}|x_{1})]$ from different pretraining stages (x-axis).
Figure 4: The relationship between y-axis $\mathcal{I}(x_{0},x_{1})+\mathcal{H}(x_{1})-\mathcal{H}(x_{0})$ and accuracy (x-axis) during fine-tuning.
4.3 Advantages
Compared to previous approaches, our method offers several key advantages:
Model-Agnostic Flexibility
Our approach is agnostic to the choice of neural network architecture for the generative model — it can be a U-Net, Transformer, or any other model. The latent variable $Z$ remains stationary with respect to the data $X$ when using an ODE solver due to the static nature of continuous-time stochastic flow models, as observed by song2020score. This allows different architectures to encode the data while yielding the same latent variable $Z$. In contrast, methods like DDAE [xiang2023denoising] are model-dependent, with the latent variable $Z$ tied to specific network architectures and derived from intermediate activations, making them less flexible and non-stationary.
Infinite-Layer Expressiveness
Our method leverages the infinite-layer structure of continuous-time stochastic flow models [Chen2018], providing high expressiveness with relatively small parameter sizes. This structure underpins the success of modern generative models like Stable Diffusion 3 [rombach2022high] and DALL·E 3 [betker2023improving]. By using a reversible process, our method allows for the simultaneous training of discriminative and generative models, offering state-of-the-art performance with this enhanced capacity.
Robustness and Generalizability
Features extracted by reversing the generative process exhibit stability along the reverse trajectory, enabling the use of features from any point $t\in[0,1]$ for downstream tasks, not just the latent variable $x_{1}$. These features are robust to different discretization schemes (e.g., Euler, Runge-Kutta) and various time steps (ranging from $10^{3}$ to $10^{1}$). Additionally, our method generalizes well across datasets, meaning that a powerful pretrained model can be fine-tuned further for improved performance on new tasks.
In summary, our approach bridges the gap between generative and discriminative models, offering a simple yet effective way to leverage pretrained generative models for unsupervised representation learning, with the ability to fine-tune for downstream supervised tasks.
5 Experiments
We conducted experiments to evaluate our method across three key aspects. First, we analyzed the effectiveness of pretraining and finetuning strategies to identify critical components of our approach. Second, we benchmarked our method against state-of-the-art generative and classical classification methods on CIFAR-10 [Krizhevsky2009LearningML], Tiny-ImageNet [Le2015TinyIV], and ImageNet [Russakovsky2014ImageNetLS]. Finally, we performed ablation studies to assess the impact of different design choices—such as architecture and optimization techniques—on downstream classification performance.
5.1 Settings
For we use three diffusion and flow models, GVP, ICFM, and OTCFM. For pretraining, we followed the protocol from [nichol2021improved], using the Adam optimizer [kingma2014adam] with a fixed learning rate of $1\times 10^{-4}$ for 1,200 epochs on the respective training sets. For downstream image classification, we adopted the configuration from [liu2021swin], using AdamW for 200 epochs with a cosine decay learning rate schedule and a 5-epoch linear warm-up. The training used a batch size of 128 and an initial learning rate of 0.001. To ensure a controlled setting, we excluded advanced regularization techniques like mixup [zhang2017mixup] and cutmix [yun2019cutmix], relying on simple data augmentation. Further details, including hyperparameters and computation efficiency, are provided in Secs. B.1 and B.2 .
5.2 Verification Analysis
We aim to address the following four questions.
5.2.1 Better Pretraining Lead to Better Finetuning?
To explore whether better pretraining improves finetuning, we compared the classification accuracies of models finetuned from different pretraining epochs on CIFAR-10 (Fig. 4). We also computed the mutual information during pretraining, which increases monotonically, as shown in Fig. 4, and analyzed the relationship between accuracy and mutual information during finetuning in Fig. 4.
Models without pretraining achieved approximately $73.5\%$ accuracy. In contrast, the more pretraining a model received, the higher its mutual information and classification accuracy, indicating that stronger generative capability leads to better finetuning performance. During finetuning, we observed a decline in $\mathcal{I}(x_{0},x_{1})+\mathcal{H}(x_{1})-\mathcal{H}(x_{0})$ for both the training and validation sets, suggesting that the model filters out unnecessary features to improve downstream classification, which can also be observed in Fig. 5 as some visual representations are kept while others are discarded. In summary, sufficient pretraining is crucial for optimal performance in inverse generative classification.
5.2.2 What is a Reasonable Finetuning Strategy?
Figure 5: Reverse generation from $x_{0}$ to $x_{1}$ using fully pretrained and fine-tuned models. See Sec. B.4 for details.
Figure 6: Accuracy on CIFAR-10 and Tiny-ImageNet when fine-tuned from different time steps along the generation trajectory.
We first examine whether a frozen flow model provides meaningful features. For this, we use features $x_{t}$ along the sampling trajectory from $x_{0}$ to $x_{1}$ under two conditions: (1) freezing the generative model’s parameters (light lines) and (2) end-to-end training of both the generative model and the classification head (dark lines). As shown in Fig. 6, the performance gap is substantial—47.10% on CIFAR-10 and 58.04% on Tiny-ImageNet—demonstrating the critical importance of updating the generative model’s parameters during finetuning.
Additionally, on CIFAR-10, initiating finetuning from later stages of the sampling trajectory ($x_{1/4}$ to $x_{1}$) results in progressively better performance. On TinyImageNet, optimal results are achieved when finetuning starts from $x_{1/2}$ to $x_{1}$. This difference likely stems from the complexity of the datasets: more complex datasets require stronger feature extraction, necessitating longer trajectory lengths for effective finetuning. Further experiments exploring optimal finetuning strategies are detailed in the ablation studies.
5.2.3 Does the Model Serve as a Continuous Feature Extractor?
Residual networks, such as ResNet-50 [he2016deep], achieve feature extraction by constructing discrete-time methods through a series of composite transformations: $\mathbf{h}_{t+1}=\mathbf{h}_{t}+f(\mathbf{h}_{t},\theta_{t})$. In contrast, we construct the feature extractor using the ODE specified by the neural network, namely $\frac{d\mathbf{h}(t)}{dt}=f(\mathbf{h}(t),t,\theta)$. Tabs. 1 and 2 demonstrate that although we trained using a discrete sample length of $t_{\text{span}}=20$, we can evaluate the model using any reasonable inference steps. Furthermore, feature extraction is not limited to a specific training point (such as the midpoint); extracting features within its vicinity (within 20%) still achieves high performance on downstream tasks. These findings validate our method can serve as a continuous feature extractor.
Inference Steps 20 100 500 800 1000
PRG-GVP Acc. (%) 97.25 97.28 97.30 97.31 97.31
PRG-ICFM Acc. (%) 97.32 97.33 97.34 97.34 97.34
PRG-OTCFM Acc. (%) 97.42 97.43 97.43 97.44 97.44
Table 1: The checkpoint, fine-tuned at the midpoint of a trajectory with $t_{\text{span}}=20$, is evaluated on CIFAR-10 using the midpoints of trajectories with varying inference steps (from 20 to 1000).
Offset from Midpoint % -20 -10 10 20
PRG-GVP Acc. (%) $97.02$ $97.17$ $97.12$ $97.08$
PRG-ICFM Acc. (%) $97.23$ $97.30$ $97.20$ $97.10$
PRG-OTCFM Acc. (%) $97.31$ $97.32$ $97.26$ $97.13$
Table 2: Accuracy at the midpoints of trajectories with varying offsets for $t_{\text{span}}=1000$ (use the same checkpoint from Tab. 1).
Method Param. (M) Acc. (%)
Discriminative methods
WideResNet-$28$-$10$ [zagoruyko2016wide] $36$ $96.3$
ResNeXt-$29$-$16*64$d [xie2017aggregated] $68$ $96.4$
Generative methods
GLOW [fetaya2019understanding] N/A $84.0$
Energy model [duvenaud2020your] N/A $92.9$
SBGC [zimmermann2021score] N/A $95.0$
HybViT [yang2022your] $43$ $96.0$
DDAE [xiang2023denoising] $36$ $97.2$
[…]
PRG-GVP-S $42$ $97.35$
PRG-ICFM-S $42$ $97.59$
PRG-OTCFM-S $42$ $97.65$
Table 3: Comparison on the CIFAR-10 dataset with various algorithms. All results are reported from their original papers.
Method Param. (M) Acc. (%)
Discriminative methods
WideResNet-$28$-$10$ [zagoruyko2016wide] $36$ $69.3$
Generative methods
HybViT [yang2022your] $43$ $56.7$
DDAE [xiang2023denoising] $40$ $69.4$
[…]
PRG-GVP-S $42$ $70.98$
PRG-ICFM-S $42$ $71.12$
PRG-OTCFM-S $42$ $71.33$
Table 4: Comparison on the Tiny-ImageNet dataset with various algorithms. All results are reported from their original papers.
Supplementary Material
Appendix A Diffusion process definition
Full definition of diffusion process that studied in this paper is a Generalized VP-SDE (GVP) using triangular function as coefficients [Albergo2023]:
$\alpha_{t}=\cos\left(\frac{1}{2}\pi t\right),$ (A1)
$\sigma_{t}=\sin\left(\frac{1}{2}\pi t\right).$
Appendix B Additional Implementation Details
B.1 Training Process Details
Tabs. B1 and B2 list the hyperparameters used for both pretraining and finetuning across the CIFAR-10, Tiny-ImageNet, and ImageNet datasets.
Dataset CIFAR-10 Tiny-ImageNet ImageNet
GPU 8$\times$×A100 8$\times$×A100 8$\times$×A100
Optimizer Adam Adam Adam
LR base 1e-4 1e-4 1e-4
Epochs 1000 1000 2000
Batch Size 256 128 128
Table B1: Experimental settings across datasets for pretraining.
Dataset CIFAR-10 Tiny-ImageNet ImageNet
GPU 4$\times$×A100 8$\times$×A100 32$\times$×A100
Optimizer AdamW AdamW AdamW
Eps 1e-8 1e-8 1e-8
Betas (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
LR base 1.25e-4 1.25e-4 1.25e-4
Weight Decay 0.05 0.05 0.05
Scheduler CosineLR CosineLR CosineLR
Warmup LR base 1.25e-7 1.25e-7 1.25e-7
Min LR base 1.25e-6 1.25e-6 1.25e-6
Epochs 200 200 200
Warmup Epochs 5 10 10
Image Size 32 64 64
Batch Size 256 128 128
T Span 20 32 64
Solver Euler Euler Euler
Table B2: Experimental settings across datasets for finetuning.
To enhance the reproducibility of results across various multi-stage and multi-GPU experiments, we calculate the learning rate using Eq. B1.
LR $=\text{LR}_{\text{base}}\times\frac{\text{num processes}\times\text{Batch Sizee}}{512}$ (B1)
Warmup LR $=\text{Warmup LR}_{\text{base}}\times\frac{\text{num processes}\times\text{Batch Size}}{512}$
[…]
B.2 Evaluation of Training Efficiency
Most research experiments, including the main experiments and ablation studies, are completed within 2 hours to 3 days. To demonstrate the training efficiency, Tab. B3 report the per-epoch runtime for each dataset used in our experiments.
$t_{\text{span}}/t_{\text{cutoff}}$ CIFAR-10 TinyImageNet ImageNet*
[…]
Table B3: Time per epoch during the training process.* means that the model is trained on $4*8$ A100 GPUs.$t_{\text{span}}$ represents the total sampling length, while $t_{\text{cutoff}}$ indicates the point along the trajectory where fine-tuning begins. For example, $20/10$ means a trajectory spanning $20$ steps from $x_{0}$ to $x_{1}$, with fine-tuning starting from the midpoint of the trajectory.
B.3 Details of Out-of-Distribution Experiments
Tiny-ImageNet-C
Method Clean Average Corruption-$5$
Adversarial Training
PGD [madry2017towards] $51.08$ $33.46$ $\downarrow 17.62$ $24.00$ $\downarrow 27.08$
PLAT [kireev2022effectiveness] $51.29$ $37.92$ $\downarrow 13.37$ $29.05$ $\downarrow 22.24$
Noise Injection
RSE [liu2018towards] $53.74$ $27.99$ $\downarrow 25.75$ $18.92$ $\downarrow 34.82$
ENResNet [wang2020enresnet] $49.26$ $25.83$ $\downarrow 23.43$ $19.01$ $\downarrow 30.25$
Data Augment
AugMix [hendrycks2019augmix] $52.82$ $37.74$ $\downarrow 15.08$ $28.66$ $\downarrow 24.16$
AutoAug [cubuk2019autoaugment] $52.63$ $35.14$ $\downarrow 17.49$ $25.36$ $\downarrow 27.27$
Generative Methods
PDE+ [yuan2024pde] $53.72$ $39.41$ $\downarrow 14.31$ $30.32$ $\downarrow 23.40$
[…]
Table B4: (OOD: extrapolation datasets) Performance on Tiny-ImageNet-C.Averge represents the accuracy across all corruption levels, with corruption severity ranging from $1$ to $5$.
There is no direct correspondence between the test images of Tiny ImageNet and Tiny ImageNet-C, and the images in Tiny ImageNet-C do not overlap with the training images of Tiny ImageNet. We report the comparative results on Tiny ImageNet-C in Tab. B4.
Figure B1: Reverse Generation Process on the Swiss Roll Dataset.
Figure B2: Reverse Generation Process on the TinyImageNet val set.
B.4 Reverse Generation Process
Fig. B1 illustrates the reverse generation process from $x_{0}$ to $x_{1}$ after finetuning. Furthermore, Figs. B2 and B3 present the reverse generation results on the TinyImageNet dataset after pretraining and finetuning, respectively. Finally, Fig. B4 demonstrates the reverse process before and after applying fog corruption to the images.
Appendix C Ablation Studies
C.1 Loss Type
Tab. C1 shows that label smoothing reduces overconfident predictions, improves model calibration, and enhances robustness.
Dataset LabelSmooth Loss Cross-Entropy Loss
CIFAR-$10$ $97.59$ $96.18$
TinyImageNet $71.12$ $70.15$
Table C1: Comparison of LabelSmooth Loss and Cross-Entropy Loss on different datasets.
C.2 The Number of Timeteps
Our findings in Tab. C2 show that longer time spans generally lead to better accuracy. On CIFAR-10 with the ICFM flow model, a $t$-span of 10 achieves accuracy comparable to the best result at $t=100$. In contrast, TinyImageNet requires a $t$-span of 15 to achieve similar performance.
T-span $\mathbf{2}$ $\mathbf{5}$ $\mathbf{10}$ $\mathbf{50}$ $\mathbf{100}$
GVP CIFAR-10 $30.54$ $90.23$ $93.26$ $97.50$ $97.55$
GVP Tiny ImageNet $5.06$ $48.95$ $53.24$ $71.05$ $71.18$
ICFM CIFAR-10 $31.18$ $92.35$ $97.02$ $97.60$ $97.61$
ICFM Tiny ImageNet $6.01$ $60.06$ $65.16$ $71.20$ $71.58$
Table C2: Comparison of Model Performance Over Varying Time Spans.
Figure B3: Reverse Generation Process on the TinyImageNet train set.The first row represents the fully pretrained reverse generative process, the second row shows the reverse generative process after extensive finetuning.
Figure B4: Reverse Generation Process on the TinyImageNet-C Dataset. The first row represents the fully pretrained reverse generative process, the second row shows the reverse process after extensive finetuning, and the third row illustrates the reverse generative process under fog corruption.